# Bounds and Bugs: The Limits of Symmetry Metrics to Detect Partisan Gerrymandering

### v1

@@ 4.2 Short Burst Analysis
Figure 15: Short burst results for Massachusetts’s State House map and 2018 US Senate election data, searching for maps with as many Democratic-won districts as possible. Declination not included because it is frequently undefined. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Note that Partisan Bias and Mean-Median values do not increase as number of districts won increases, as the other metrics do.
@@
Figure 16: Short burst results for Massachusetts’s State House map and 2018 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Declination not included because it is frequently undefined. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Note that Partisan Bias and Mean-Median values do not decrease as number of districts won increases, as the other metrics do.
@@
Here we can see that the range of values for all metrics except MM and PB is higher when the number of Democratic-won districts is higher. And conversely, the range of values for all metrics except MM and PB is lower when the number of Republic-won districts is higher. For the Mean-Median Difference and Partisan Bias, the range of values tends to be similar, regardless of the number of districts won by a particular party. These are the results found in [30], except that [30] did not consider the Partisan Bias, which is included here. Thus, we can similarly conclude that the Partisan Bias is not useful in determining extreme maps, just like the Mean-Median Difference.
@@
We note that these results (indicating that the symmetry measures of Partisan Bias and Mean-Median Difference values are not useful in flagging extreme maps) are similar to those in [10], although here we see those results on 18 maps, rather than the three maps studied in [10]. The results for the other maps can be found in Appendix C.
5 Discussion
The results from our Short Burst analyses, such as those in Figures 15 and 16, show that the values of the Partisan Bias and Mean-Median difference do not indicate whether or not a map is an extreme map. The values do not generally increase as the number of Democratic-won districts increase (or decrease as the number of Republican-won districts decrease) as the values of the other metrics do. This observation, along with our discussion of “districts won” being the foundation of gerrymandering outlined in Section 1.2, leads us to suggest that the Partisan Bias and Mean-Median Difference are not accurate metrics at determining partisan gerrymandering.
But these results also invite the corresponding question: Why are the other metrics more consistent with districts won? What is it about the other metrics that leads them to increase when Democratic-won districts increase, and decrease when Democratic-won districts decrease? The answer is clear with the Efficiency Gap, as it is simply a function in the vote-share, seat-share, and a turnout ratio $\rho$ [34]. Specifically, if
$\rho=\frac{\text{average turnout in districts party }A\text{ lost}}{\text{average turnout in districts party }A\text{ won}}$
then
$EG=\left(S-\frac{1}{2}\right)-2\left(V-\frac{1}{2}\right)+\frac{S(1-S)(1-\rho)}{S(1-\rho)+\rho}$
Thus, for a neutral ensemble of maps using fixed partisan data (and thus fixed $V$), as $S$ goes up, EG will also increase [^9] Yes, $\rho$ will also change, but its change is relatively small. This is why the bands of potential EG values for different $S$ are narrow..
For the Geography and Election Outcome (GEO) metric, the argument is more intuitive. Recall that the Democratic GEO score of a map is the number of additional districts that could have potentially been made competitive for the Democratic party, and the Republican GEO score is the number of additional districts that could have potentially been made competitive for the Republican party [6]. The GEO score ratio used in the Short Burst analysis (used because it has values between -1 and 1, similar to other metrics, and thus we could display the green “reasonable band” as with the other metrics) is:
Rep GEO scoreDem GEO scoretotal number of $\frac{\text{Rep GEO score}-\text{Dem GEO score}}{\text{total number of districts}}$
For a neutral ensemble of maps, evaluated using the same partisan data (and thus fixed $V$), consider two maps with different Democratic seat shares. One can imagine that the map $M_{\text{Dem higher}}$ with higher Democratic seat share can be constructed from the map $M_{\text{Dem lower}}$ with lower Democratic seat share by adjusting $M_{\text{Dem lower}}$. Using the ideas behind the GEO metric, this may be the result of changing some of the districts that could have become competitive for the Democratic party into districts won by the Democratic party. If this is the case, the GEO score for the Democratic party in $M_{\text{Dem higher}}$ would be lower. Conversely, those districts that are now won by the Democratic party in $M_{\text{Dem higher}}$ could become competitive for the Republican party, suggesting that the GEO score for the Republican party in $M_{\text{Dem higher}}$ would be higher. Thus, this suggests that the GEO ratio for $M_{\text{Dem higher}}$ would be larger than the GEO ratio for $M_{\text{Dem lower}}$, which would explain why the GEO ratio tends to increase as the Democratic-won districts increases in the neutral ensemble.
To explain why the Declination also tends to increase as Democratic-won districts increases, we can look to the proof of Theorem 1 in [36]. What the proof shows is the following: suppose $M_{\text{Dem lower}}$ and $M_{\text{Dem higher}}$ are maps with the same partisan data (and thus same $V$) such that $M_{\text{Dem lower}}$ has lower Democratic seat share than $M_{\text{Dem higher}}$. Let $\overline{V_{\ell}}(M)$ be the average Democratic vote share in districts that Democrats lots in map M and $\overline{V_{w}}(M)$ be the average Democratic vote share in districts that Democrats won in map $M$. Then if we have both
$\overline{V_{\ell}}(M_{\text{Dem higher}})$ $<\overline{V_{\ell}}(M_{\text{Dem lower}})$ (1)
$\overline{V_{w}}(M_{\text{Dem higher}})$ $<\overline{V_{w}}(M_{\text{Dem lower}})$ (2)
then
$\delta(M_{\text{Dem lower}})<\delta(M_{\text{Dem higher}})$
where $\delta$ is the Declination. That is, if the average vote shares in districts that Democrats lost tends to get lower as $S$ increases, and if the average vote shares in districts that Democrats won tends to get lower as $S$ increases, then the Declination tends to increase as $S$ increases. Intuitively, one can understand how this might happen through packing and cracking, although it certainly may not always occur.
For example, consider two different elections in a state with 5 districts. The first election results in two districts having vote shares of 30% and three districts having vote share 80%. The second election results in one district having vote share $\frac{7}{15}\approx 46.667\%$ and four districts having vote share $\frac{19}{30}\approx 63.333\%$. The vote shares and declination values of both elections are the same ($V=60\%,\delta=0$). However the seat shares are different. Thus, for this pair of elections, the declination cannot detect the seat share change. (We note that not both of (1) and (2) can be true for this pair of elections.)
However, while equations (1) and (2) may not always be true in constructed data, they may be frequently true in real data. To assess this empirically, we created histograms of $\overline{V_{\ell}}$ and $\overline{V_{w}}$ for the neutral ensembles we created for each map and each election. Typical results can be seen in Figure 17.
Figure 17: Histograms on the left are the average vote shares in districts Democrats lost, color coded by the number of districts won by Democrats. Histograms on the right are the average vote shares in districts Democrats won, color coded by the number of districts won by Democrats. First row is for Michigan congressional, second row is Texas Congressional, third row is Pennsylvania congressional. All are using election data from the 2016 Presidential election. All show that the histogram values tend to move higher as the number of districts won by Democrats decreases.
In Figure 17, we can see that the values of $\overline{V_{\ell}}$ do tend to increase as $S$ increases, and the values of $\overline{V_{w}}$ also tend to increase as $S$ increases. This gives a suggestion as to why the Declination tends to increase as $S$ increases.
@@ 6 Conclusion
In this article, for each possible $(V,S)$ pair that can be achieved with turnout equal in each district, we have calculated the possible range of values of $MM$ and $PB$ for constructed election data with vote-share $V$ and seat-share $S$. We have seen that this range of values behaves in ways that are not what one would naturally expect of a metric intended to detect partisan gerrymandering.
As a result of our calculations, we have completed a long-missing study of the Mean-Median Difference and Partisan Bias, which had been completed for other metrics: showing which seats-votes pairs $(V,S)$ have constructed election data giving $MM=0$ or $PB=0$. We have shown that these $(V,S)$ pairs are the same for both metrics, and cover a wide swath of the $[0,1]\times[0,1]$ square. This region expands when we allow the ratio $C$ of maximum turnout to minimum turnout to be larger than 1; so much so that the region fills the $[0,1]\times[0,1]$ square in the limit as $C\to\infty$. This suggests that the Mean-Median Difference and Partisan Bias metrics are potentially less reliable when detecting gerrymanders.
Indeed, we empirically show that, although the signs of the Mean-Median Difference and Partisan Bias must be the same, the values of the two metrics can be quite different. The largest theoretical difference is 0.5, and a difference of 0.33 is seen in South Carolina’s Congressional map (with several other maps having a difference of 0.2 or more in absolute value).
The study in [30] showed that the Mean-Median Difference can give misleading results even when evaluated on a neutral ensemble, and we show here that the same is true of the Partisan Bias.
The study in [30] showed that the range of values of the Mean-Median Difference is similar for both extreme maps (maps with an extreme number of districts won by a particular party) and non-extreme maps, and we show similar results here for the Partisan Bias. We discuss reasons why the Efficiency Gap, Declination, and GEO metric are better at tracking “districts won” than the Mean-Median Difference and Partisan Bias, and back up the reasoning for the Declination with empirical studies.
Warnings have been raised about symmetry metrics such as the Mean-Median Difference and Partisan Bias in [10], stating “For those who do want to constrain the most extreme partisan outcomes that line drawing can secure, these investigations should serve as a strong caution regarding the use of partisan symmetry metrics, whether in the plan adoption stage or in plan evaluation after subsequent elections have been conducted.” The additional theoretical and empirical findings in this article give even more evidence that partisan symmetry metric values fail to accurately detect when a map has an extreme number of districts won by a particular party. When used in a neutral ensemble, an outlier Mean-Median Difference of Partisan Bias may or may not correspond to an outlier in number of districts won by a given party. We similarly strongly caution against using these metrics at any stage. They simply are not reliable at flagging extreme maps, even within a neutral ensemble.
@@ 7 Acknowledgements
The author would like to thank Thomas Ratliff for sharing the code he wrote to make the beautiful images for the Short Burst analysis results, most of which can be seen in Appendix C. The author would also like to thank Moon Duchin, whose comments, questions, and suggestions have greatly improved this article.
Appendix A Proofs of Theorems 1, 2, and 3
@@
Firstly note that, since we assume that $S>\frac{1}{2}$, the median vote share must be at least as large as $\frac{1}{2}$. For $V\leq\frac{1}{2}$, we can achieve a median vote share of $\frac{1}{2}$ by constructing $S$ districts to have vote share $\frac{1}{2}$ (considered winning districts) and choosing the losing districts to have vote shares so that $\frac{1}{n}\sum_{i=1}^{n}V_{i}=V$, since we must have $S\leq 2V$. This would give vote share $V$, seat share $S$, and the smallest possible median of $\frac{1}{2}$.
@@
Appendix B Neutral ensemble Images
Figure 18: Neutral ensembles for Massachusett’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric. Declination value of 2 used for maps where Declination is undefined.
Figure 19: Neutral ensembles for Michigan’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 20: Neutral ensembles for Michigan’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 21: Neutral ensembles for Michigan’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 22: Neutral ensembles for Oregon’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 23: Neutral ensembles for Oregon’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 24: Neutral ensembles for Oregon’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 25: Neutral ensembles for Oklahoma’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric. Declination not included because it is not defined for a majority of maps.
Figure 26: Neutral ensembles for Oklahoma’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 27: Neutral ensembles for Oklahoma’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 28: Neutral ensembles for Pennsylvania’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 29: Neutral ensembles for Pennsylvania’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 30: Neutral ensembles for Texas’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 31: Neutral ensembles for Texas’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Appendix C Short Burst Images
Figure 32: Short burst results for Massachusetts’s Congressional map and 2018 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for nearly all maps.
Figure 33: Short burst results for Massachusetts’s Congressional map and 2018 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for nearly all maps.
Figure 34: Short burst results for Massachusetts’s State Senate map and 2018 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for several maps.
Figure 35: Short burst results for Massachusetts’s State Senate map and 2018 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for several maps.
Figure 36: Short burst results for Michigan’s Congressional map and 2016 Presidential election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 37: Short burst results for Michigan’s Congressional map and 2016 Presidential election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 38: Short burst results for Michigan’s State Senate map and 2016 Presidential election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 39: Short burst results for Michigan’s State Senate map and 2016 Presidential election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 40: Short burst results for Michigan’s State House map and 2016 Presidential election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 41: Short burst results for Michigan’s State House map and 2016 Presidential election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 42: Short burst results for Oklahoma’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for a majority of the maps.
Figure 43: Short burst results for Oklahoma’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for a majority of the maps.
Figure 44: Short burst results for Oklahoma’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for a majority of the maps.
Figure 45: Short burst results for Oklahoma’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for a majority of the maps.
Figure 46: Short burst results for Oklahoma’s State House map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 47: Short burst results for Oklahoma’s State House map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 48: Short burst results for Oregon’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 49: Short burst results for Oregon’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 50: Short burst results for Oregon’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 51: Short burst results for Oregon’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 52: Short burst results for Oregon’s State House map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 53: Short burst results for Oregon’s State House map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.

### v3

@@ 5.2 Results of Empirical Study
Figure 10: Empirical results for Massachusetts’s State House map and 2018 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
@@
Figure 11: Empirical results for Massachusetts’s State House map and 2018 US Senate election data, searching for maps with as many Republican-won districts as possible. Declination not included because it is frequently undefined. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
@@
In Figures 10 and 11, we can see that the range of values for the Efficiency Gap and GEO Ratio is higher when the number of Democratic-won districts is higher. And conversely, the range of values for the Efficiency Gap and GEO Ratio is lower when the number of Republic-won districts is higher. To reiterate, this is what would be expected for a metric which detects an extreme number of districts won: the metric has a range of values that is more extreme on maps with a more extreme number of districts won. In contrast, for the Mean-Median Difference and Partisan Bias, the range of values of those metrics is similar, regardless of the number of districts won by a particular party [^21] These are the results found in [45], except that [45] did not consider the Partisan Bias, which is included here.. Additionally, these ranges of values does not seem to have an upward trend on the maps with an extreme number of Democratic districts nor a downward trend on the maps with an extreme number of Republican districts.
These results, that extreme values of the Mean-Median Difference and Partisan Bias do not correlate with extreme number of districts won, are seen on every map we tested. We will explore a few more examples; the reader interested in seeing the others can find them in Appendix B.
While the previous example (Massachusets State House, in Figures 10 and 11) corresponds to a very Democratic-leaning state and a map with many districts (160), the next example is for a very Republican-leaning state with few districts (5): Oklahoma’s congressional map. The results for that map are in Figures 12 and 13.
Figure 12: Empirical results for Oklahoma’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 13: Empirical results for Oklahoma’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Again, in Figures 12 and 13, we see that the range of values for the Mean-Median Difference and Partisan Bias are not more extreme on maps with a more extreme number of districts won by the party in question. In fact, while the mean values of the Efficiency Gap and GEO ratio increase as the number of districts won by Democrats increases (and the mean values decrease as the number of districts won by Republicans increases), as expected, those mean values do the opposite for the Mean-Median Difference and Partisan Bias. That is, the “average” value of the Mean-Median Difference (Partisan Bias) on maps with an extreme number of Democratic districts suggests that they are less gerrymandered towards Democrats than maps with a fewer number of Democratic districts won [^22] And the “average” value of the Mean-Median Difference (Partisan Bias) on maps with an extreme number of Republican districts suggests that they are less gerrymandered towards Republicans than maps with a fewer number of Republican districts won..
The third and final example result that we give in this section is for a “purple” state (one not clearly Democratic or Republican leaning): Pennsylvania. In Figures 14 and 15, we see the results for Pennsylvania’s congressional map. Again, we see that the most extreme values of the Mean-Median Difference and Partisan Bias are achieved on many maps, not just those with the most extreme number of districts won by a particular party.
Figure 14: Empirical results for Pennsylvania’s Congressional map and 2016 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 15: Empirical results for Pennsylvania’s Congressional map and 2016 US Senate election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figures 10, 11, 12, 13, 14, 15, and the other Figures in Appendix B give clear evidence that extreme values of the Mean-Median Difference and Partisan Bias do not correspond to maps with an extreme number of districts won. That is, while we saw in Section 4 that the values these two metrics are able to take on do not become more extreme when a more extreme number of districts is won, here we see that, even on real maps/data, the values that these two metrics do take on cannot detect extreme number of districts won.
@@
We note that these results (indicating that the symmetry metricss of Partisan Bias and Mean-Median Difference values are not useful in flagging extreme maps) are similar to those in [15], although here we see those results on 18 maps, rather than the three maps studied in [15]. Thus, while the results of [15] could potentially have been dismissed as “flukes” and maps specifically chosen because of some kind of unique property, our results robustly show that on every single map of the 18 maps we considered, the Mean-Median Difference and Partisan Bias do not detect an extreme number of districts won. We encourage the reader to view the results for the other maps, which can be found in Appendix B.
@@ 6 Conclusion
In this article, we considered the relationship between possible values of the ubiquitous Mean-Median Difference and Partisan Bias metrics and the number of seats won under a fixed districting plan as a proxy for detecting gerrymandering. The results are contained in two studies: a theoretical study (the “bounds” study) and an empirical study (the “bugs” study). In our theoretical study, we saw how the values the MM and PB could take on were different for different vote-share, seat-share pairs. We saw that the values these metrics could take on are not, in general, more extreme when a more extreme number of districts is won. We also saw how these metrics can take on values that run counter to the public’s expectations, and to the ways they are frequently used and described. In our empirical study, we considered 18 different maps: 3 maps for each of 6 states with varying political geographies. These 18 maps also had a wildly varying number of districts (ranging from 5 to 203). In each of these maps, we saw that more extreme values of the Mean-Median Difference and Partisan Bias did not correspond to maps with more extreme number of districts won.
Given the evidence we present, we hope that experts and the general public take extreme care when attempting to use these metrics in the future. We also hope that public-facing apps, experts, and academics will provide clear descriptions of these metrics and their limitations in order to more accurately describe what they can and cannot do. We are not the first to give this warning [15], though our paper provides significant additional evidence that this warning should be heeded. Extreme values of these metrics simply do not correspond to an extreme number of districts won, and thus these metrics should not be solely used for that purpose.
@@ 7 Acknowledgements
The authors would like to thank Thomas Ratliff for sharing the code he wrote to make the beautiful images for the Empirical analysis results, most of which can be seen in Appendix B. The authors would also like to thank Moon Duchin, whose comments, questions, and suggestions have greatly improved this article. Finally, the authors would like to thank the reviewers for their careful reading and thoughtful comments.
This material is based upon work supported by the National Science Foundation under Grant No. DMS-1928930, while the authors were in residence at the Mathematical Sciences Research Institute in Berkeley, California, during the semester/year of Fall 2023.
Appendix A Proofs of Theorems 1, 3, and 5
@@
Firstly note that, since we assume that $S>\frac{1}{2}$, the median vote share must be at least as large as $\frac{1}{2}$. For $V\leq\frac{1}{2}$, we can achieve a median vote share of $\frac{1}{2}$ by constructing $S$ of the districts to have vote share $\frac{1}{2}$ (considered winning districts) and choosing the losing districts to have vote shares so that $\frac{1}{n}\sum_{i=1}^{n}V_{i}=V$, since we must have $S\leq 2V$. This would give vote share $V$, seat share $S$, and the smallest possible median of $\frac{1}{2}$.
@@
Appendix B Images from Empirical Analysis